OpenStudy (anonymous):

Use Cramer's Rule to solve the system or to determine that the system is inconsistent or contains dependedent equations. 3x+2y=2 6x+5y=1

OpenStudy (amistre64):

1 2 1 6 1 5 cover a column take the determinant of whats left, and repeat that process 3 times to get all the needed numbers

OpenStudy (amistre64):

D(c1,c2) = Dx D(c2,c3) = Dy D(c1,c3) = D x = Dx/D y = Dy/D

OpenStudy (amistre64):

grrr.... 3 2 2 6 1 5 thatll do it , i loathe this tine little laptop

OpenStudy (amistre64):

Dy=3(1) - 6(2) = -9 Dx=2(5) - 1(2) = 8 D = 3(5) - 6(2) = 3 y = -9/3 x = 8/3 and dbl chk to make sure :)

OpenStudy (anonymous):

thanks